Constants| Periodic Table Part A A 6000 kg lunar lander is in orbit 90 km above the surface of the moon. It needs to move out to a 300 km -high orbit in order to link up with the mother ship that will take the astronauts home. How much work must the thrusters do? Express your answer with the appropriate units. 4.97 1015 J Submit Previous Answers Request Answer X Incorrect; Try Again; 7 attempts remaining
Answer #1

total energy = half of the potential energy

wordone = final total energy - initial total energy

= -(1/2)G*M_moon*m/(Rm+h2) - ((-(1/2)G*M_moon*m/(Rm+h1))

=(1/2) G*M_moon*m*( 1/(Rm+h1) - 1/(Rm+h2))

M_moon = 7.35*10^22 kg

Rm = 1737.4 km

W = (1/2)6.67*10^-11*7.35*10^22*6000*(1/(1737.4*10^3+90*10^3) - 1/(1737.4*10^3+300*10^3)

W = 8.30*10^8 J